Opened 7 years ago

Last modified 6 years ago

#464 new flight processing

GB12/05, flight day 259/2012, Monks Wood — at Version 16

Reported by: knpa Owned by:
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by edfi)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/uk/GB12_05-2012_259_Monks_Wood

Data arrived from ARSF via SATA disk on 24/09/2012

GB12/05 Scientific objective: unknown

GB12/05 Priority: a4m

PI: Ross Hill

Sensors:

Camera (13/12/12)
Eagle
Hawk
LiDAR
FW LiDAR

Change History (16)

comment:1 Changed 7 years ago by knpa

  • Summary changed from GB12/05, flight day 249/2012, Monks Wood to GB12/05, flight day 259/2012, Monks Wood

comment:2 Changed 7 years ago by besm

Navigation processed.

comment:3 Changed 7 years ago by besm

Eagle and hawk raw files are listed as SWIR259a-*, but the flight has been renamed from 259a to 259. Hyperspectral processing doesn't work. I'm renaming them to SWIR259-* now.

comment:4 Changed 7 years ago by anhi

RCD Delivery Created

Last edited 7 years ago by anhi (previous) (diff)

comment:5 Changed 7 years ago by edfi

Beginning Lidar

comment:6 Changed 7 years ago by anhi

  • Description modified (diff)

comment:7 Changed 7 years ago by anhi

Delivered Camera to Ross Hill sent on 13th December.

comment:8 Changed 7 years ago by besm

Hyperspectral delivery proceeding without vectors. Delivery created and readme is ongoing.

comment:9 Changed 7 years ago by besm

Readme created. Ready to delivery check.

Please make sure to check all files are in the right place and appropriately named, as the make delivery script may have had a silent problem.

comment:10 Changed 7 years ago by edfi

Started D/C of Hyperspectral

comment:11 Changed 7 years ago by edfi

Only three bands processed needs re-processing

comment:12 Changed 7 years ago by besm

Starting reprocessing.

comment:13 Changed 7 years ago by besm

Finished reprocessing. Ready to re-delivery check.

comment:14 Changed 7 years ago by edfi

Started Hyperspectral delivery check

comment:15 Changed 7 years ago by edfi

Hyperspectral check complete was missing log sheet which has now been added. Everything is fine. Currently zipping mapped files

comment:16 Changed 7 years ago by edfi

  • Description modified (diff)
Note: See TracTickets for help on using tickets.